OpenStudy (anonymous):

My answer is B What is the solution to the rational equation 3/2 - 1/x - 5 = -2/2x - 10? x = -5/2

OpenStudy (anonymous):

\[\frac{3}{2} - \frac{1}{x - 5} = \frac{-2}{2x - 10}\] is my guess, but i could be wrong

OpenStudy (anonymous):

@InsanelyChaotic does that look right?

OpenStudy (anonymous):

It looks right to me

OpenStudy (anonymous):

very good job

hartnn (hartnn):

then that doesn't have a solution! how you got x=-5/2 ??

OpenStudy (anonymous):

because its in the equation and it's negative!!!....oooops

hartnn (hartnn):

-5/2 is in the equation ? i don't see it :P

OpenStudy (anonymous):

there is no solution, but we can try and solve it and see what happens

OpenStudy (anonymous):

subtract on the left and get \[\frac{3(x-5)-2}{2x-10}\] or \[\frac{3x-15-2}{2x-5}=\frac{3x-17}{2x-10}\] then since the denominators are the same on the left and right, you get \[3x-17=-2\] solve this and get \[3x-17=-2\] \[3x=15\] \[x=5\]

OpenStudy (anonymous):

but \(x\) cannot be \(5\) because if \(x\) were \(5\) the denominator would be \(5-5=0\) and you cannot have a zero in the denominator

OpenStudy (anonymous):

great info guys....thanks for the explanations and not just flat out answers

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i would give a medal to both if i could! xo

hartnn (hartnn):

i am not here for medals... so you can give to him if you want :)